OpenStudy (anonymous):

There are two numbers whose sum is 30 and whose product is 221. This number differ by (A) 4 (B) 3 (C) 2 (D) 1

OpenStudy (anonymous):

A

OpenStudy (anonymous):

First problem: Call the numbers x and y Then: x + y = 30 xy = 221 y = 30 - x (subtract x from each side of first equation) x(30 - x) = 221 (substitute 30-x for y in second equation) 30x - x^2 = 221 distributive property of multiplication over addition -x^2 + 30x = 221 re-arrange terms on left side -x^2 + 30x - 221 = 0 subtract 221 from each side Feed the coefficients into the quadratic formula to find that x = 13 or 17, so the two numbers differ by 4.

OpenStudy (anonymous):

oh...okay.